OpenStudy (anonymous):

(2+x) dy/dx=3y How do I solve this with separation of variables?

OpenStudy (anonymous):

(2+x)dy/dx = 3y (1/3y)dy = 1/(2+x) dx (1/3)ln(y) = ln(x+2) + c

OpenStudy (agent0smith):

\[(2+x) \frac { dy} {dx}=3y\] can be rearranged to: \[\frac{ dy }{ 3y } = \frac{ dx }{(2+x) }\]

OpenStudy (anonymous):

(1/3)ln(y) = ln(x+2) + c ln(y) = 3*ln(x+2) + c ; c*3 is just another constant c y = e^(3*ln(x+2) + c) y = e^ln((x+2)^3) *e^c y = (x+2)^3 *e^c // e^c is just another constant c y = (x+2)^3 *c

OpenStudy (anonymous):

Don't you need to integrate somewhere?

OpenStudy (anonymous):

// sorry i didn't show it explicitly (1/3y)dy = 1/(2+x) dx // here i integrated. (1/3)ln(y) = ln(x+2) + c

OpenStudy (anonymous):

Oh ok. Thanks. Makes more sense. I kind of caught it, but I wanted to make sure.

OpenStudy (agent0smith):

\[\int\limits \frac{ dy }{ 3y } = \int\limits \frac{ dx }{(2+x) }\] do you know how to integrate from here: \[\frac{ 1 }{ 3 } \int\limits \frac{ 1 }{ y } dy = \int\limits \frac{ 1}{(2+x) } dx\]